The diameter of the base of a cylinder is 22 cm. The height of the cylinder is 16 cm. Find the volume of the cylinder. Round to the nearest hundredth.
Answer by lwsshak3(11628)

The diameter of the base of a cylinder is 22 cm. The height of the cylinder is 16 cm. Find the volume of the cylinder. Round to the nearest hundredth.
**
Volume of a cylinder=area of base*height
Area of base=πd^2/4=π*22^2/4=121π
volume=16*121π=1936π≈6082.123 cm^3